OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (hba):

sorry dont know

OpenStudy (australopithecus):

this is a line so you have the general formula y = mx + b MEMORIZE THIS where m = slope b = y-intercept

OpenStudy (australopithecus):

calculate the slope first since you have two points you can easily do this with the formula m = (y2-y1)/(x2-x1)

OpenStudy (australopithecus):

\[m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (australopithecus):

sub both points into the equation to find the slope (x, y) (-5, 3) (1, -3)

OpenStudy (anonymous):

i got -6/6

OpenStudy (australopithecus):

so you got a slope of -6/6 = -1

OpenStudy (australopithecus):

so sub that into the equation of the line which you should memorize it. y = (-1)x + b

OpenStudy (australopithecus):

all you have to do is sub a point into the equation of the line to solve for b

OpenStudy (australopithecus):

that will give you the y intercept

OpenStudy (anonymous):

but its asking for the x-intercept?

OpenStudy (australopithecus):

yes with the equation for the line you can find the x intercept

OpenStudy (australopithecus):

you are finding the equation of the line from these two points with that equation you can figure out what the x-intercept is by setting y = 0

OpenStudy (australopithecus):

you can also graph the line with it

OpenStudy (australopithecus):

did you solve for b?

OpenStudy (anonymous):

i got -2 for b

OpenStudy (australopithecus):

alright so you have the equation y = -x - 2 now you know your y intercept is at point (0,-2) now we need to find the x-intercept so (x,0) so set y = 0 0 = -x - 2 solve for x and you will have your x-intercept and the question will be finished

OpenStudy (anonymous):

i got -2 is that correct?

OpenStudy (australopithecus):

yes you can look at the graph here http://www.wolframalpha.com/input/?i=y+%3D+-x+-+2+

OpenStudy (anonymous):

thanksssss :D

OpenStudy (australopithecus):

Although the correct way to answer it is y-intercept (0,-2) x-intercept (-2,0)

OpenStudy (anonymous):

maybe use y=mx+b